### v4

# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

Abstract
Computational efficiency and non-adversarial robustness are critical factors in process modeling and optimization for real-world engineering applications. Yet, conventional neural networks often fall short in addressing both simultaneously, or even separately. Drawing insights from natural physical systems and existing literature, it is known theoretically that an input convex architecture will enhance computational efficiency, while a Lipschitz-constrained architecture will bolster non-adversarial robustness. However, integrating both properties into one model is a nontrivial task, as enforcing one property may compromise the other one. Therefore, in this work, we develop a novel network architecture, termed Input Convex Lipschitz Recurrent Neural Networks, that inherits the strengths of both convexity and Lipschitz continuity. This model is explicitly designed for fast and robust optimization-based tasks, which outperforms existing recurrent units in terms of computational efficiency and non-adversarial robustness. Additionally, we have successfully implemented this model in various practical engineering applications, such as optimization of chemical processes and real-world solar irradiance prediction for Solar PV system planning at LHT Holdings in Singapore. Source code is available at https://github.com/killingbear999/ICLRNN.
1 Introduction
Neural networks provide a powerful tool to process modeling, an important engineering problem in various manufacturing industries such as chemical, pharmaceutical, and energy systems. This work draws inspiration from nature-inspired design to develop a fast and robust solution that addresses two crucial challenges: computational efficiency and non-adversarial robustness, for modeling of real-world nonlinear systems. Specifically, nerual network models have been developed to capture complex nonlinear dynamics of industrial manufacturing systems, and the models can be further used for process monitoring, predictive maintenance, optimization, and real-time control. However, while neural network-based optimization can optimize process performance such as energy consumption and process profits, based on the prediction of neural networks, computational efficiency remains a big challenge for its real-time implementation. For example, model predictive control (MPC), an advanced optimization-based control technique, has been widely used to manage a process while complying with a set of constraints. Over the past decades, neural network-based MPC (NN-MPC) has received increasing attention, where a neural network that models system dynamics is incorporated into the design of MPC optimization problems [1, 2, 3, 4, 5, 6, 7, 8]. However, general optimization problems (e.g., MPC) built on conventional neural networks often fall short of meeting the requirement of computational efficiency. Nonetheless, efforts have been made to mitigate these limitations. For example, the computational efficiency of NN-MPC is enhanced by implementing an input convex architecture in neural networks [9]. Additionally, non-adversarial robustness of NN-MPC is improved by developing Lipschitz-constrained neural networks (LNNs) [10].
In real-world applications, computational efficiency plays a pivotal role due to the imperative need for real-time decision-making, especially in critical engineering processes such as chemical process operations [9]. We define computational efficiency as the speed at which computational tasks are completed within a given execution timeframe. An effective approach to achieve this efficiency is through the transformation from non-convex to convex structures. Convexity is a ubiquitous characteristic observed in various physical systems such as the potential energy function of a mass on a spring. Building upon nature-inspired design and extending its application to neural networks, Input Convex Neural Networks (ICNNs) is a class of neural network architectures intentionally crafted to maintain convexity in their output with respect to the input. Leveraging these models could be particularly advantageous in various engineering problems, notably in applications such as MPC, due to their inherent benefits especially in optimization. This concept was first introduced as Feedforward Neural Networks (FNN) [11], and subsequently extended to Recurrent Neural Networks (RNN) [12], and Long Short-Term Memory (LSTM) [9].
Additionally, in real-world engineering applications, non-adversarial robustness plays a critical role due to the prevalent noise present in most real-world data, which significantly hampers neural network performance. Our definition of robustness aligns closely with the principles outlined by [10]. Given the inherent noise in real-world process data, our goal is to improve neural networks by effectively learning from noisy training data for comprehensive end-to-end applications. Leveraging LNNs offers a potential solution to the robustness challenge. This concept was first introduced as FNN [13], and subsequently extended to RNN [14], and Convolutional Neural Networks (CNN) [15]. However, at this stage, the development of neural networks that address the above two challenges simultaneously remains an open question. It is important to note that the integration of ICNNs and FNNs presents a non-trivial challenge due to their interplay in the way that enforcing one property in the design of neural networks may compromise the other property.
Motivated by the above considerations, in this work, we introduce an Input Convex Lipschitz RNN (ICLRNN) that amalgamates the strengths of ICNNs and LNNs. This approach aims to achieve the concurrent resolution of computational efficiency and non-adversarial robustness for real-world engineering applications. Specifically, we validate the efficacy of ICLRNN across various engineering problems that include neural network-based optimization for chemical processes and time series forecasting in a real-world Solar Photovoltaic (PV) system in Singapore. Performance assessment hinges on modeling accuracy, computational efficiency (i.e., evaluated by runtime and floating-point operations), and non-adversarial robustness. Our objective is to achieve equilibrium across these three dimensions as they hold equal importance in real-world engineering scenarios.
In summary, Section 2 offers an overview of modern high-performing recurrent units, ICNNs and LNNs. Section 3 presents the proposed ICLRNN, substantiating its theoretical attributes as input convex and Lipschitz continuous. Section 4 illustrates that ICLRNN surpasses state-of-the-art recurrent units in various engineering tasks, including optimization and control of chemical processes using MPC and real-time solar irradiance forecasting on Solar PV systems. Section 5 delves into the potential limitations of ICLRNN uncovered through experiments conducted on the Long Range Arena benchmark [16].
2 Background
2.1 Notations
[…]
2.2 High-performing RNNs
In recent years, several high-performing recurrent units have emerged, including UniCORNN [17], rank-coding based RNNs [18], and Linear Recurrent Units [19]. However, these variants primarily emphasize enhancing accuracy, especially for very long time-dependent sequences modeling, which often at the expense of model simplicity and computational efficiency. Moreover, in many engineering applications that we are currently working on with industry partners, accuracy alone does not dictate the suitability of a model. In fact, from an engineering point of view, simplicity and robustness are also important in addition to accuracy, as they ensure the model can be easily adapted and incorporated into existing industrial operational systems, and can be trusted in an uncertain environment with noise and disturbances. Therefore, the goal of this work is to develop a network architecture based on simple RNNs that strikes a balance between these factors while maintaining a satisfactory level of accuracy, rather than investigating the novel designs of RNNs for improved accuracy only.
2.3 Input Convex RNN
The initiative behind ICNNs is to leverage the power of neural networks, specifically designed to maintain convexity within their decision boundaries [11]. ICNNs aim to combine the strengths of neural networks in modeling complex data with the advantages of convex optimization, which ensures the convergence to a global optimal solution. They are particularly attractive for control and optimization problems, where achieving globally optimal solutions is essential. By maintaining convexity, ICNNs help ensure that the optimization problems associated with neural networks (e.g., neural network-based optimization) remain tractable, thus addressing some of the challenges of non-convexity and the potential for suboptimal local solutions in traditional neural networks. A foundational work, known as Input Convex RNN (ICRNN) [12], serves as one of the baselines.
2.4 Lipschitz RNN
By adhering to the definition of Lipschitz continuity for a function $f$, where an $L$-Lipschitz continuous $f$ ensures that any minor perturbation to the input results in an output change of at most $L$ times the magnitude of that perturbation. Therefore, constraining neural networks to maintain Lipschitz continuity significantly fortifies their resilience against input perturbations. A pivotal work, namely the Lipschitz RNN (LRNN) [14], stands as one of the baselines in this paper.
3 Input Convex Lipschitz Recurrent Neural Networks
3.1 Convexity and Lipschitz Continuity of ICLRNN
We propose a novel network architecture, termed ICLRNN, that possesses both Lipschitz continuity and input convexity properties. Specifically, the output of ICLRNN is defined as follows:
$h_{t}=g_{1}(W_{t}^{x}x_{t}+W_{t}^{h}h_{t-1}+b_{t}^{h})$ (1a)
$y_{t}=g_{2}(W_{t}^{y}h_{t}+b_{t}^{y})$ (1b)
where all $W_{i}$ are constrained to be non-negative with singular values small and bounded by $\mathcal{O}(1)$, and all $g_{i}$ are constrained to be convex, non-decreasing, and Lipschitz continuous. Similar to LRNN and ICRNN, UniCORNN and Linear Recurrent Units involve high Floating Point Operations, and do not provide a theoretical guarantee on non-adversarial robustness and computational efficiency on convex neural-network-based optimization problem. These aspects motivated us to retain the standard RNN architecture while imposing input convexity and Lipschitz constraints on the weights, since many engineering applications can achieve satisfactory modeling accuracy even using relatively simple NN structures. However, how to improve the robustness and computational efficiency of these NN models remains an important question.
The output of a neural network inherits input convexity and Lipschitz continuity if and only if every hidden state possesses these properties. Unlike [12] and [14], our ICLRNN adheres to this foundational principle by imposing constraints on the weights and activation functions instead of introducing supplementary variables to the original RNN architecture. This approach minimizes the demand of computational resources. The following proposition demonstrates that following our proposed method of imposing an input convex constraint subsequent to a Lipschitz constraint does not compromise the latter, following the proposed design of ICLRNNs. The proofs of the following Propositions and Theorems can be found in Appendix A.
We first establish Proposition 1 to demonstrate that the imposition of an Input Convex constraint subsequent to a Lipschitz constraint does not compromise the Lipschtiz property to a significant extent, allowing both to harmoniously coexist.
Proposition 1
Let $A$ be an $m\times n$ matrix with its largest singular value at most 1. If all negative elements in matrix $A$ are replaced by 0 to obtain a non-negative matrix $B$, then the largest singular value of matrix $B$ remains small and bounded by $\sqrt{r(A)}$, and further bounded by $\sqrt{min(m,n)}$.
For ICLRNN, we first enforce the spectral constraints to ensure that the largest singular values of the weight matrix is at most one [13]. The spectral constraint is implemented in two steps, which is similar to [15]. Firstly, we use spectral normalization as proposed by [20], where the spectral norm is reduced to at most 1 by iteratively evaluating the largest singular value with the power iteration algorithm, which is a method used to find the dominant eigenvector of a square matrix using an iterative numerical technique that repeatedly multiplies the matrix by a vector and normalizes the result until convergence. Secondly, we apply the Björck algorithm, which is an iterative algorithm for computing the best estimate of an orthogonal matrix (i.e., iteratively refining the initial estimate of the orthogonal matrix by minimizing the error between the estimate and the true orthogonal matrix, eventually converging to a stable solution), to increase other singular values to at most one.
Next, we enforce the non-negative constraint by performing weight clipping to set all negative values in the weights to 0. The theoretical upper bound after the integration of the input convex constraint with the Lipschitz constraint is $\sqrt{r(W)}$ for one single time step as shown in Proposition 1. It is essential to note that the upper bound (i.e., $\|A\|=\sigma_{max}(A)\leq\|A\|_{F}$) is not precisely tight [21], and thus the upper bound developed in Proposition 1 is not tight. Through experimentation on random large matrices and the experiments in Section 4, we have observed that the Lipschitz constant is normally bounded within the order of $\mathcal{O}(1)$.
Moreover, ICLRNN uses ReLU as the activation function for hidden states instead of the GroupSort function proposed in [13]. This decision aims to maintain convexity within the model architecture. Although GroupSort exhibits gradient norm preservation and higher expressive power [10], its non-convex nature contrasts with our primary objectives. Next, by combining Proposition 1, Proposition 2, and Proposition 3, Theorem 1 indicates the convexity and Lipschitz continuity of ICLRNN.
Proposition 2
The Lipschitz constant of an ICLRNN is $\mathcal{O}(1)$ if and only if all weights $W_{i}$ and all activation functions $g_{i}$ have a Lipschitz constant bounded by $\mathcal{O}(1)$.
Proposition 3
The output $y$ of an ICLRNN is input convex if and only if all $W_{i}$ are constrained to be non-negative and all $g_{i}$ are constrained to be convex and non-decreasing.
Theorem 1
The output $y$ of an L-layer Input Convex Lipschitz Recurrent Neural Network is a convex, non-decreasing, and Lipschitz continuous function with respect to the input $x$, where $x$ is in a convex feasible space $\mathcal{X}$, if and only if the following constraints on weights $W_{i}$ and activation functions $g_{i}$ are satisfied simultaneously: (1) All $W_{i}$ are constrained to be non-negative with singular values bounded by $\mathcal{O}(1)$; (2) All $g_{i}$ are constrained to be convex, non-decreasing, and Lipschitz continuous (e.g., ReLU, Linear, Softmax).
To demonstrate the benefits of ICLRNN in neural-network-based optimization tasks, we proceed by constructing an MPC framework, one of the most widely-used control methods in chemical industries that regulates the process states to the desired set point by solving optimization problems in real time. The goal is to show that a non-convex NN-based optimization problem can be converted into a convex optimization problem through the utilization of ICNNs.
3.2 ICLRNN for a finite-horizon convex MPC
3.2.1 Class of systems
We consider a category of systems that can be expressed through the ordinary differential equations (ODEs) of the following form:
$\dot{x}=F(x,u)$ (2)
where $u\in\mathbb{R}^{m}$ denotes the control action and $x\in\mathbb{R}^{n}$ denotes the state vector. The function $F:X\times U\to\mathbb{R}^{n}$ is continuously differentiable, where $X\subset\mathbb{R}^{n}$ and $U\subset\mathbb{R}^{m}$ are connected and compact subsets that enclose an open neighborhood around the origin respectively. In this context, we maintain the assumption throughout this task that $F(0,0)=0$, ensuring that the origin $(x,u)=(0,0)$ serves as an equilibrium point. As it might be impractical to obtain first-principles models for intricate real-world systems, our objective is to construct a neural network to model the nonlinear system specified by Eq. (2), embed it into optimization problems, and maintain computational efficiency and non-adversarial robustness simultaneously.
3.2.2 MPC formulation
The optimization problem of the Lyapunov-based MPC (LMPC) that incorporates a ICLRNN as its predictive model and uses a Lyapunov function as stability constraints, is expressed as follows:
[…]
$s.t.\ \dot{\tilde{x}}(t)=F_{nn}(\tilde{x}(t),u(t))$ (3b)
$u(t)\in U,\ \forall t\in[t_{k},t_{k+N})$ (3c)
$\tilde{x}(t_{k})=x(t_{k})$ (3d)
$V({\tilde{x}}(t))<V({x}(t_{k})),\forall t\in[t_{k},t_{k+N})$ (3e)
where $S(\triangle)$ denotes the set of piecewise constant functions with period $\triangle$, $\tilde{x}$ denotes the predicted state trajectory, and $N$ denotes the number of sampling periods in the prediction horizon. The objective function $\mathcal{L}$ outlined in Eq. (3a) integrates a cost function $J$ that depends on both the control actions $u$ and the system states $x$. The system dynamic function $F_{nn}(\tilde{x}(t),u(t))$ expressed in Eq. (3b) is parameterized by an RNN (e.g., the proposed ICLRNN in this work). Eq. (3c) encapsulates the constraint function $U$, delineating feasible control actions. Eq. (3d) establishes the initial condition $\tilde{x}(t_{k})$ in Eq. (3b), referring to the state measurement at $t=t_{k}$. Lastly, Eq. (3e) represents the Lyapunov-based constraint $V$, ensuring closed-loop stability within LMPC by mandating a decrease in the value of $V(x)$ (i.e., Lyapunov functions $V(x)$ are analogous to energy functions in some sense that the system is stable at the lowest energy point where $V(x)=0$ at $x=0$).
It is important to highlight that a convex neural network-based LMPC remains convex even when making multi-step ahead predictions, where the prediction horizon is greater than one [9]. This statement holds true if we incorporate an inherently input convex neural network (e.g., ICLRNN) to the LMPC, as described in Eq. (3b), and if we guarantee that the cost function $J$ in Eq. (3a) is input convex throughout the task.
4 Empirical evaluation
In this section, we assess the performance of the proposed ICLRNN by benchmarking it against other state-of-the-art recurrent units. We aim to exhibit the computational efficiency and non-adversarial robustness of ICLRNN across various engineering challenges. Incorporating real-world data from our industrial partner can illustrate the practical applicability of our proposed method. Unlike abstract benchmark tasks, our real-world scenario aligns more closely with engineering contexts. Demonstrating effectiveness in a practical setting not only enhances the credibility of our approach but also serves commercial interests and addresses stakeholder needs.
Drawing from the above considerations, we meticulously choose two case studies to showcase the computational efficiency and non-adversarial robustness of ICLRNN. These include implementing an NN-MPC for a chemical reactor and predicting solar irradiance for Solar PV systems using real-time data from LHT Holdings plants in Singapore. It is worth noting that ablation studies are conducted to ascertain the optimal hyperparameter values for all subsequent tasks. Moreover, given that our research encompasses both Lipschitz-constrained NNs and Input Convex NNs, we meticulously select the most suitable approach from these categories and ensure consistency by employing a unified set of baseline methods across both tasks. Our primary aim, however, is to showcase the superiority of our proposed method over any alternatives within these two distinct families. Details on the computing resources are provided in Appendix D.
4.1 Application to real-time optimization of continuous stirred tank reactor
MPC is a real-time optimization-based control scheme, where the first element of the optimal input trajectory will be executed in the system over the next sampling period, and the MPC optimization problem will be resolved at the subsequent sampling period iteratively until convergence. Note that in general, solving NN-MPC is time-consuming, since it is a non-convex optimization problem due to the nonlinearity and nonconvexity of neural network models. However, if we integrate an ICNN into a convex MPC problem, the MPC problem will remain convex. In the next subsection, we demonstrate the improvement of the non-adversarial robustness in chemical process modeling and computational efficiency of NN-MPC using the proposed ICLRNN.
4.1.1 Nonlinear chemical process modeling
We implement MPC to a nonisothermal and well-mixed continuous stirred tank reactor (CSTR) featuring an irreversible second-order exothermic reaction. This reaction facilitates the conversion of reactant A into product B. The CSTR incorporates a heating jacket responsible for either supplying or extracting heat at a rate $Q$. The dynamic model of the CSTR is characterized by the ensuing material and energy balance equations:
$\frac{dC_{A}}{dt}=\frac{F}{V}(C_{A0}-C_{A})-k_{0}e^{\frac{-E}{RT}}C_{A}^{2}$ (4a)
$\frac{dT}{dt}=\frac{F}{V}(T_{0}-T)+\frac{-\triangle H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}^{2}+\frac{Q}{\rho_{L}C_{p}V}$ (4b)
where $C_{A}$ denotes the concentration of reactant A, $F$ denotes the volumetric flow rate, $V$ denotes the volume of the reactant, $C_{A0}$ denotes the inlet concentration of reactant A, $k_{0}$ denotes the pre-exponential constant, $E$ denotes the activation energy, $R$ denotes the ideal gas constant, $T$ denotes the temperature, $T_{0}$ denotes the inlet temperature, $\triangle H$ denotes the enthalpy of reaction, $\rho_{L}$ denotes the constant density of the reactant, $C_{p}$ denotes the heat capacity, and $Q$ denotes the heat input rate. The detailed parameter values are listed in Table 2 in Appendix B.
Moreover, $x^{T}=[C_{A}-C_{As},T-T_{s}]$ denotes the system states, where $T-T_{s}$ and $C_{A}-C_{As}$ represent the deviations in the temperature of the reactor and the concentration of A from their respective steady-state values. $u^{T}=[\triangle C_{A0},\triangle Q]$ denotes the control actions, where $\triangle Q=Q-Q_{s}$ and $\triangle C_{A0}=C_{A0}-C_{A0_{s}}$ denote the manipulated inputs within this system, representing the alterations in the heat input rate and the inlet concentration of reactant A, respectively. The objective of the MPC of Eq. (3) involves operating the CSTR at $(C_{As},T_{s})$, which is the unstable equilibrium point, by manipulating $\triangle Q$ and $\triangle C_{A0}$ respectively. In summary, the inputs consist of $[T_{t}-T_{s},C_{A_{t}}-C_{As},Q_{t},C_{{A0}_{t}}]$ at the current time step $t$. The outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps (i.e., representing a one-to-many sequence problem).
We conduct open-loop simulations for the CSTR of Eq. (4) using explicit Euler method. These simulations aim to collect data that mimics the real-world data for training purposes (i.e., as it is difficult to obtain real-world chemical plant data). All models undergo training and evaluation using uniform configurations, incorporating the mean-squared loss (MSE), the Adam Optimizer, and the ReLU activation function. The MPC reaches the convergence state when $|x_{1}|=|C_{A}-C_{As}|<0.1~{}kmol/m^{3}$ and $|x_{2}|=|T-T_{s}|<3~{}K$ simultaneously. The initial conditions are chosen within the stability region, which is defined as $01060x_{1}^{2}+44x_{1}x_{2}+0.52x_{2}^{2}-372=$ 0 (i.e., an ellipse in state space). For a comprehensive evaluation of model performance across various hypothesis space sizes and initial conditions, detailed outcomes and discussions are presented in Appendix B.
4.1.2 Non-NN methods performance
In addition to NN-based methods, we investigate various data-driven models to provide a more comprehensive analysis, including non-NN methods such as the linear state-space model and the nonlinear autoregressive network with exogenous inputs (NARX), which have been widely used for modeling nonlinear dynamic systems for decades. Both the linear state-space model (i.e., testing MSE of $1.17\times 10^{-3}$) and the NARX model (i.e., testing MSE of $6.37\times 10^{-4}$) fail to outperform even a standard RNN (i.e., testing MSE of $9.77\times 10^{-5}$).
4.1.3 Performance analysis for CSTR modeling
Based on our experiments, we observed that a testing MSE on the order of $10^{-3}$ is indicative of satisfactory performance, rendering the model suitable for applications such as NN-MPC (i.e., NN-MPC can converge). To compare the computational efficiency in modeling the system of Eq. (4), we quantify their Floating Point Operations per Second (FLOPs). Table 1 illustrates that ICLRNN has the lowest FLOPs, thereby outperforming or matching state-of-the-art methods. It is crucial to emphasize that the FLOPs presented in Table 1 are calculated for models that share identical architectures, including the same number of layers and neurons.
Table 1: Model complexity in terms of FLOPs.
Model FLOPs (CSTR) FLOPs (Solar PV)
RNN 406,548 399,362
LSTM 1,597,460 1,596,418
ICRNN 1,204,233 1,195,010
LRNN 2,505,748 2,498,562
ICLRNN (ours) 406,548 399,362
To scrutinize the non-adversarial robustness of ICLRNN in modeling the system dynamics in Eq. (4), we assess the model’s sensitivity with respect to test accuracy when exposed to a series of perturbed inputs influenced by Gaussian noise. As illustrated in Fig. 1(a), the findings, averaging across 5 trials, reveal that ICLRNN consistently outperforms or matches the performance of state-of-the-art models across varying levels of noise. These findings indicate ICLRNN’s non-adversarial robustness across diverse hypothesis space sizes with respect to different Gaussian noise levels. Despite LRNN having a relatively modest testing MSE against different Gaussian noise levels, its high FLOPs indicate substantial computational demands, which are about 6 times greater than ICLRNN. This computational intensity stands as a significant drawback despite its competitive performance. Notably, ICLRNN exhibits superior computational efficiency, as evidenced by its notably lower number of FLOPs compared to other models. Moreover, ICRNN encounters challenges when the hypothesis space expands, succumbing to the issue of the exploding gradient problem, which makes it difficult to train.
4.1.4 Performance analysis for CSTR-based LMPC
Next, incorporating an ICNN into the aforementioned LMPC in Eq. (3) transforms it into a convex optimization problem. Subsequently, PyIpopt, as a Python version of IPOPT [22], is used to address the LMPC problem. The integration time step is set at $h_{c}=1\times 10^{-4}~{}hr$, while the sampling period remains $\triangle=5\times 10^{-3}~{}hr$. The design of the Lyapunov function, denoted as $V(x)=x^{T}Px$, involves configuring the positive definite matrix $P$ as $\big{[}\begin{smallmatrix}1060&22\\ 22&0.52\end{smallmatrix}\big{]}$, which ensures the convexity of the LMPC. We execute LMPC for a fixed timeframe for various models and repeat with various initial conditions to evaluate the performance of different NN-based LMPC. Fig. 1(b) shows one of the initial conditions, from which it is evident that the ICLRNN-based LMPC converges, whereas other NN-based LMPC models do not. The fastest convergence speed highlights the superior computational efficiency of the ICLRNN-based LMPC. Moreover, it is important to highlight that ICLRNN-based LMPC consistently demonstrates the fastest convergence speed across various initial conditions, thereby surpassing the state-of-the-art methodologies. Performance with respect to other initial conditions are discussed in Appendix B.
(a) Model performance w.r.t different degree of Gaussian noise for CSTR.
(b) LMPC converging path in a fixed timeframe with an initial condition at $[-1.5~{}kmol/m^{3},70~{}K]$.
Figure 1: Performance of CSTR-based LMPC.
4.2 Application to real-world solar irradiance prediction
In this task, we aim to forecast solar irradiance using real-time data for LHT Holdings, a wood pallet manufacturing industry based in Singapore. Before the installation of the Solar photovoltaic (PV) system, LHT Holdings relied solely on the main utility grid to fulfill its energy needs. The Solar PV system was successfully installed by 10 Degree Solar in late 2022 (i.e., a detailed manufacturing pipeline of LHT Holdings and the solar PV system are available in Appendix C). This transition to solar energy was primarily driven by the aim of minimizing manufacturing costs. Three significant uncertainties were identified within the industry, namely solar irradiance affecting the Solar PV system’s efficiency, the dynamic real-time pricing of energy sourced from the main utility grid, and the fluctuating energy demand stemming from unexpected customer orders.
Currently, the industrial facility primarily relies on Solar PV as its main energy source, with the main utility grid serving as a backup. Excess solar energy is stored for later use. Accurate solar irradiance forecasting is crucial for efficient energy planning and cost reduction. Using sensors installed by the Solar Energy Research Institute of Singapore (SERIS) and Singapore Institute of Manufacturing Technology (SIMTech), including irradiance, humidity, wind, and temperature sensors, alongside data similar to that used by [23], we predict minute-by-minute solar irradiance to optimize energy dispatch. In this time-series forecasting task, all models undergo training and evaluation using identical settings. These settings encompass MSE as the loss function, the Adam Optimizer, the ReLU activation function. For the training and validation phases, we use data spanning from January 14, 2023, to December 16, 2023. Subsequently, data from December 17, 2023, to January 1, 2024, is exclusively reserved for testing and evaluating the models’ predictive performance.
Furthermore, from an engineering perspective, the training and implementation of neural network models are often constrained by limited computational resources, which are typically not as ample as those in computer science. Consequently, engineering applications necessitate relatively simple models capable of effectively handling challenges such as data noise and convexity. To ensure the selection of the most suitable model, it is essential to conduct comparisons under consistent settings (i.e., the same model complexity). This approach ensures fairness, as we evaluate not only accuracy but also other critical factors such as computational efficiency, which can be quantified by FLOPs. The comprehensive discussion of results is accessible in Appendix C.
4.2.1 Performance analysis for solar irradiance forecasting
We leverage FLOPs as a measure to assess the computational efficiency of models. As depicted in Table 1, it is evident that among all the models, ICLRNN demonstrates the most efficient computational performance as a result of its lowest FLOPs. Consequently, this superior efficiency positions ICLRNN as the top-performing model when considering computational resources.
Given the real-world context of this task, introducing artificial noise for evaluating non-adversarial robustness becomes unnecessary as the inherent noise within real-world data suffices. Notably, our observations reveal that ICRNN encounters challenges with the exploding gradient problem when the hypothesis space size is large (i.e., the number of hidden neurons per layer is more than 128), thereby hindering its ability to effectively model complex system dynamics. While RNN and LSTM exhibit a limited capacity to capture system dynamics, they struggle notably when faced with sudden changes in solar irradiance, leading to less accurate predictions.
In Fig. 2(a), we present results from the best state-of-the-art model, LRNN, alongside our model’s outcomes specifically for December 28, 2023. The LRNN, constrained by Lipschitz continuity, exhibits significantly improved performance compared to the RNN and LSTM models. However, it still encounters challenges in accurately predicting solar irradiance during sudden fluctuations, and its excessively high FLOPs signify its computational intensity (i.e., 2,498,562 for LRNN compared to 399,362 for ICLRNN). Moreover, slight fluctuations in the LRNN predictions occur during periods of minimal solar irradiance at the start and end of the day. Highlighted within Fig. 2(b), our ICLRNN outperforms LRNN and other state-of-the-art models, particularly in accurately predicting sudden changes in solar irradiance, which highlights the superior predictive capability of ICLRNN, showcasing its effectiveness in capturing and predicting nuanced fluctuations within the solar irradiance data. Notably, these observations persist across all trials and are not specific solely to December 28, 2023.
(a) LRNN performances on 2023-12-28 for solar irradiance prediction.
(b) ICLRNN performances on 2023-12-28 for solar irradiance prediction.
Figure 2: Performance of solar irradiance forecasting.
5 More applications and discussions
During our experiments, we observed that ICRNN faces challenges related to the exploding gradient problem, prompting us to discuss this issue and its relevance to the exploding and vanishing gradients phenomenon. Fortunately, as a beneficial side-effect of addressing our aforementioned objectives, our proposed method exhibits reduced susceptibility to the exploding gradient problem to a certain extent, attributable to its incorporation of Lipschitz constraints. Subsequently, we further investigated the potential of ICLRNN by delving into the Long Range Arena benchmark [16], encompassing tasks involving long sequence modeling such as byte-level text sentiment classification and pixel sequence image classification. Both ICRNN and our ICLRNN exhibit limitations in tasks that require the modeling of very long time-dependent sequences, such as the byte-level text classification task, where both achieve an accuracy of approximately 50% in this binary sentiment classification task. This limitation stems from the challenge of the exploding gradient problem and lack of representative power due to non-negative weights, which both methods encounter when attempting to model extremely lengthy sequences, and we provide a theoretical explanation as follows.
While the Lipschitz constraint may alleviate the exploding gradient problem by bounding the weights, integrating input convexity introduces a new challenge by raising this upper bound. Applying the Lipschitz constraint first to reduce the spectral norm to 1, followed by imposing the input convex constraint, yields a neural network no longer 1-Lipschitz constrained but rather $n\sqrt{r(W)}$-constrained, where $n$ represents the number of time-steps. Consequently, in modeling very long time-dependent sequences, the Input Convex structure exacerbates the exploding gradient problem, rendering it unmitigated even with Layer Normalization. For readers interested in tackling the challenges of the exploding and vanishing gradient problem within RNNs, we recommend exploring several noteworthy and comprehensive works [24, 25, 26]. We acknowledge this limitation as inherent to our proposed method, thereby constraining its applicability. However, despite this drawback, we opt to retain the Input Convex structure, as it demonstrates efficacy in many engineering applications where the length of time-series data is not as long as those in Long Range Arena benchmark. For example, in the first example where we develop a model for real-time optimization of CSTR, the model is to predict one sampling period only, since the controller is typically implemented in a feedback manner with the true measurements of states received at every time step. In the second case study where we develop a model for real-world solar irradiance prediction, again, the prediction horizon is relatively short (e.g., an hour or a few hours ahead) since this model is used for real-time energy planning. Therefore, there are many applications in engineering disciplines that focus on short-term prediction, where the proposed ICLRNN will play an important role due to its superiority in computational efficiency and robustness compared to traditional RNNs.
6 Conclusion
This work introduces an Input Convex Lipschitz RNN to improve computational efficiency and non-adversarial robustness of neural networks in various engineering applications such as real-world solar irradiance prediction, and process modeling and optimization. The ICLRNN model was demonstrated to surpass state-of-the-art recurrent units across various engineering tasks, ranging from simulated studies to real-world applications, in both computational efficiency and non-adversarial robustness.
Appendix A Proofs
The definitions of convexity and Lipschitz continuity are given as follows:
Definition 1 ([27])
A function $f:X\rightarrow\mathbb{R}$ is convex if and only if the following inequality holds $\forall~{}(x_{1},x_{2})\in X$, with $x_{1}\neq x_{2}$, and $\forall~{}\lambda\in(0,1)$:
$f(\lambda x_{1}+(1-\lambda)x_{2})\leq\lambda f(x_{1})+(1-\lambda)f(x_{2})$
and it is strictly convex if the relation $\leq$ becomes $<$.
Definition 2 ([28])
A function $f:X\rightarrow\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L$ (i.e., L-Lipschitz) if and only if the following inequality holds $\forall~{}(x_{1},x_{2})\in X$:
$\|f(x_{1})-f(x_{2})\|_{2}\leq L\|x_{1}-x_{2}\|_{2}$
The following lemmas are provided to prove Proposition 4 regarding the Lipschitz continuity of ICLRNN.
Lemma 1 ([28])
Consider a set of functions, e.g., $f:X\rightarrow\mathbb{R}$ and $g:X\rightarrow\mathbb{R}$ with Lipschitz constants $L(f)$ and $L(g)$, respectively, by taking their sum, i.e., $h=f+g$, the Lipschitz constant $L(h)$ of $h$ satisfies:
$L(h)\leq L(f)+L(g)$
Lemma 2 ([28])
Consider a set of functions, e.g., $f:X\rightarrow\mathbb{R}$ and $g:X\rightarrow\mathbb{R}$ with Lipschitz constants $L(f)$ and $L(g)$ respectively, by taking their composition, denoted as $h=f\circ g$, the Lipschitz constant $L(h)$ of the resultant function $h$ satisfies:
$L(h)\leq L(f)\times L(g)$
Lemma 3
Given a linear function with a weight $W\in\mathbb{R}^{n\times m}$, we have:
$L(W)=\|W\|=\sigma_{max}(W)$
Lemma 4 ([29, 30])
The most common activation functions such as ReLU, Sigmoid, and Tanh have a Lipschitz constant that equals 1, while Softmax has a Lipschitz constant bounded by 1.
Proposition 4
The Lipschitz constant of an ICLRNN is $\mathcal{O}(1)$ if and only if all weights $W_{i}$ and all activation functions $g_{i}$ have a Lipschitz constant bounded by $\mathcal{O}(1)$.
Proof 1
By unrolling the recurrent operations over time $t$ in Eq. (1), the output of ICLRNN can be described as a function:
$f(x)=g_{t}(W_{t}^{y}g_{t-1}^{h}(W_{t-1}^{x}x_{t-1}+W_{t-1}^{h}...g_{1}(W_{1}^{x}x_{1}+W_{1}^{h}h_{0})))$ (5)
The Lipschitz constant of an ICLRNN is upper bounded by the product of sum of the individual Lipschitz constants (i.e., Lemma 1 and Lemma 2):
$L(f)\leq L(g_{t})\times L(W_{t}^{y})\times L(g_{t-1})\times(L(W_{t-1}^{x})+L(W_{t-1}^{h})\times...L(g_{1})\times(L(W_{1}^{x})+L(W_{1}^{h})))$
If we ensure that the Lipschitz constants of all $W_{i}$ in ICLRNN are $\mathcal{O}(1)$, the ICLRNN for classification tasks with ReLU activation for hidden states and Softmax activation for output layer has a Lipschitz constant of $\mathcal{O}(1)$, while the ICLRNN for regression tasks with ReLU activation for hidden states and Linear activation for output layer has a Lipschitz constant of $\mathcal{O}(1)$ (i.e., Lemma 3 and Lemma 4).
The following lemmas are provided to prove Proposition 5 regarding the convexity of ICLRNN.
Lemma 5 ([27])
Consider a set of convex functions, e.g., $f:X\rightarrow\mathbb{R}$ and $g:X\rightarrow\mathbb{R}$, their weighted sum, i.e., $\alpha f+\beta g$, remains convex if coefficients $\alpha$ and $\beta$ are non-negative.
Lemma 6 ([27])
Consider a convex function $f:X\rightarrow\mathbb{R}$ and a convex monotone non-decreasing function $g:X\rightarrow\mathbb{R}$, the composition of $f$ with $g$, denoted as $g\circ f$, is convex and non-decreasing.
Proposition 5
The output $y$ of an ICLRNN is input convex if and only if all $W_{i}$ are constrained to be non-negative and all $g_{i}$ are constrained to be convex and non-decreasing.
Proof 2
By unrolling the recurrent operations over time $t$ in Eq. (1), the output of ICLRNN can be described as a function of Eq. (5). The output of an ICLRNN $y=f(x)$ is convex with respect to $x$ (i.e., Lemma 5 and Lemma 6).
The subsequent Proposition 6 aims to establish an upper bound on the weights for a single time step, following the integration of the input convex constraint with the Lipschitz constraint.
Proposition 6
Let $A$ be an $m\times n$ matrix with its largest singular value at most 1. If all negative elements in matrix $A$ are replaced by 0 to obtain a non-negative matrix $B$, then the largest singular value of matrix $B$ remains small and bounded by $\sqrt{r(A)}$, and further bounded by $\sqrt{min(m,n)}$.
Proof 3
We begin the proof with a commonly admitted fact:
$\|A\|=\sigma_{max}(A)\leq\|A\|_{F}=(\sum_{ij}|a_{ij}|^{2})^{\frac{1}{2}}=(\sum_{k}\sigma_{k}^{2})^{\frac{1}{2}}$ (6)
where $A$ is an $m\times n$ matrix with its largest singular value at most 1 (i.e., $\|A\|=\sigma_{max}(A)\leq 1$), $a_{ij}$ represents the elements in $A$, and the number of singular value $k$ is at most $min(m,n)$. Equality holds if and only if $A$ is a rank-one matrix or a zero matrix. Given that the rank of a matrix is precisely the number of non-zero singular values, we extend the inequality in Eq. (6) further:
$\|A\|_{F}\leq(r(A)\sigma_{max}(A)^{2})^{\frac{1}{2}}=\sqrt{r(A)}\|A\|$ (7)
where the equality holds if and only if all non-zero singular values are equal.
Let $B$ be a non-negative $m\times n$ matrix transformed by replacing all negative elements in $A$ with 0, where $b_{ij}$ represents the elements in $B$. Based on the transformation, we have the property:
$|b_{ij}|^{2}\leq|a_{ij}|^{2}$
Thus, the following inequality holds:
$\|B\|\leq\|B\|_{F}=(\sum_{ij}|b_{ij}|^{2})^{\frac{1}{2}}\leq(\sum_{ij}|a_{ij}|^{2})^{\frac{1}{2}}=\|A\|_{F}$ (8)
Based on Eq. (7) and (8) and the assumption that $\|A\|\leq 1$, we conclude:
$\|B\|\leq\|A\|_{F}\leq\sqrt{r(A)}\|A\|\leq\sqrt{r(A)}\leq\sqrt{min(m,n)}$
Appendix B MPC-based CSTR results
Table 2: Parameter values of CSTR.
$F=5~{}m^{3}/hr$ $V=1~{}m^{3}$ $R=8.314~{}kJ/kmol~{}K$
$T_{0}=300~{}K$ $C_{p}=0.231~{}kJ/kg~{}K$ $\rho_{L}=1000~{}kg/m^{3}$
$Q_{s}=0.0~{}kJ/hr$ $E=5\times 10^{4}~{}kJ/kmol$ $C_{A0_{s}}=4~{}kmol/m^{3}$
$k_{0}=8.46\times 10^{6}~{}m^{3}/kmol~{}hr$ $\Delta H=-1.15\times 10^{4}~{}kJ/kmol$
Table 1 indicates the parameter values for the CSTR. Fig. 1(a) showcases the model’s testing accuracy with respect to varying degrees of Gaussian noise applied to the input. These findings indicate ICLRNN’s non-adversarial robustness across diverse hypothesis space sizes with respect to different Gaussian noise levels. Notably, ICLRNN exhibits superior computational efficiency, as evidenced by its notably lower number of FLOPs compared to other models.
While ICRNN performs well against various degrees of Gaussian noise within smaller hypothesis spaces, it encounters challenges when the hypothesis space expands, succumbing to the issue of the exploding gradient problem. LRNN demonstrates comparable performance to ICLRNN specifically with a layer size of (512, 512). However, it is computationally demanding, characterized by substantially higher FLOPs (i.e., 9,992,212 for LRNN compared to 1,599,508 for ICLRNN). This computational intensity stands as a significant drawback despite its competitive performance.
We now delve into the control performance assessment of NN-based LMPC. To assess and compare the convergence rates of various NN-based LMPC, we conduct different NN-based LMPC under a fixed timeframe and repeat with different initial conditions. Figs. 3 and 4 provide compelling evidence that the ICLRNN-based LMPC achieves convergence, unlike other NN-based LMPC, within a fixed timeframe. This outcome distinctly indicates that ICLRNN-based LMPC exhibits the fastest convergence speed under identical settings.
In Fig. 3, the left panels show the results corresponding to the initial condition set at $[-1.5~{}kmol/m^{3},70~{}K]$. Specifically, the top left panel displays the concentration profile, the middle left panel shows the temperature, and the bottom left panel exhibits the converging path. Additionally, the right panels showcase the results from the initial condition at $[1.5~{}kmol/m^{3},-70~{}K]$, featuring the concentration profile in the top right panel, temperature profile in the middle right panel, and the converging path in the bottom right panel. Moving to Fig. 4, the left panels depict the results obtained from the initial condition of $[-1.25~{}kmol/m^{3},50~{}K]$. Here, the top left panel represents the concentration, the middle left panel illustrates the temperature, and the bottom left panel shows the converging path. Similarly, the right panels exhibit the results of the initial condition at $[1.25~{}kmol/m^{3},-50~{}K]$.
Across various initial conditions, ICLRNN-based LMPC consistently demonstrates the fastest convergence speed, thereby surpassing the state-of-the-art methodologies.
Figure 3: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.5~{}kmol/m^{3},70~{}K]$ under LMPC, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.5~{}kmol/m^{3},-70~{}K]$.
Figure 4: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.25~{}kmol/m^{3},50~{}K]$, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.25~{}kmol/m^{3},-50~{}K]$.
Appendix C Real-world solar irradiance prediction results
At present, the Solar PV system serves as the primary energy source for the industrial facility, while the main utility grid acts as a secondary energy source to supplement any deficiencies in solar energy production. Any surplus solar energy generated beyond the current requirements is efficiently stored in batteries for future utilization. Hence, in order to optimize real-time energy planning and mitigate manufacturing costs associated with energy procurement from the main utility grid, a high level of precision and accuracy in forecasting solar irradiance becomes imperative. Moreover, a simple and robust trained NN with relatively high accuracy can be seamlessly integrated into applications like energy dispatch optimization. To achieve these goals, the Solar Energy Research Institute of Singapore (SERIS) and Singapore Institute of Manufacturing Technology (SIMTech) have played pivotal roles by installing various sensors specifically designed for the Solar PV system. These sensors include irradiance sensors, humidity sensors, wind sensors, and module temperature sensors. Drawing on parameters akin to those employed by [23], we utilize minute-based average global solar irradiance, ambient humidity, module temperature, wind speed, and wind direction to predict solar irradiance in the subsequent minute.
Fig. 5 illustrates the production pipeline adopted by LHT Holdings. Meanwhile, in Fig. 6(a), a depiction of the different energy sources employed within the production process is presented. Fig. 6(b) showcases the physical setup of the Solar PV system, encompassing solar panels, inverters, batteries, meters, humidity and wind sensors, module temperature sensors, and irradiance sensors. Data is systematically recorded on a minute-by-minute basis and seamlessly uploaded online into the system.
Figure 5: LHT Holdings technical wood production pipeline.
(a) LHT Holdings energy sources for manufacturing.
(b) LHT Holdings Solar PV system.
Figure 6: Performance of solar irradiance forecasting.
Fig. 7 shows the model performance on solar irradiance prediction on December 28, 2023. The top two figures display the RNN and LSTM prediction outcomes, respectively. Both models demonstrate poor performance, failing to adjust adequately when faced with sudden changes in solar irradiance, leading to overshooting. Moreover, the middle two figures reveal that the ICRNN encounters issues related to the exploding gradient problem, particularly when the hypothesis space size becomes excessively large. The middle left figure shows the prediction performance of ICRNN with 128 hidden neurons per layer, while the middle right figure shows the prediction performance of ICRNN with 256 hidden neurons per layer. When surpassing 128 hidden neurons per layer, the ICRNN struggles to effectively capture and learn the system’s dynamics.
Moving to the bottom two figures, the LRNN and ICLRNN prediction results for the same date, December 28, 2023, are presented. The LRNN, constrained by Lipschitz continuity, exhibits significantly improved performance compared to the RNN and LSTM models. However, it still encounters challenges in accurately predicting solar irradiance during sudden fluctuations, and its excessively high FLOPs signify its computational intensity (i.e., 2,498,562 for LRNN compared to 399,362 for ICLRNN). Moreover, slight fluctuations in the LRNN predictions occur during periods of minimal solar irradiance at the start and end of the day. Notably, these observations persist across all trials and are not specific solely to December 28, 2023. It is clearly shown in Fig. 2(a) that our ICLRNN model surpasses state-of-the-art models, demonstrating superior predictive capabilities across various scenarios.
Figure 7: RNN (top left), LSTM (top right), ICRNN (middle left and middle right), LRNN (bottom left), and ICLRNN (bottom right) performance on 2023-12-28 for solar irradiance prediction.
Appendix D Computing resources
The experiments were conducted on the NVIDIA Quadro RTX A5000 GPU with 24 GB memory, but they are also compatible with CPUs equipped with at least 64 GB of RAM. A single run to train the real-world solar irradiance prediction model took about 6 hours while a single run for NN-MPC is about 20 minutes using CPU with 64 GB of RAM.